OpenStudy (anonymous):

what  are cross products and how do they relate to proportions? please help!!

OpenStudy (texaschic101):

Cross products are in problems involving proportions. We can use cross products to test whether two ratios are equal and form a true proportion. Example : 2/3 = 4/6.....now you would cross multiply (2)(6) = (4)(3)....the cross product is the result of multiplication 12 = 12....these cross products are equal, therefore, this is a true proportion.

OpenStudy (anonymous):

thank you soooo much